# Can LLMs Predict Their Own Failures? Self-Awareness via Internal Circuits

### v1

@@ 1 Introduction
Figure 1: Overview of our Gnosis self-awareness mechanism and its performance. Left: Gnosis taps hidden states and attention maps from a frozen LLM, learns to compress them into hidden/attention descriptors, and predicts a scalar correctness (hallucination) score with only $\sim$5 million extra parameters and essentially zero added inference cost. Right: Gnosis surpasses 8B Skywork reward models and a Gemini 2.5 Pro judge (hundreds-of-billions parameters) in AUROC on math reasoning, open-domain QA, and academic knowledge; scores are averaged over the tested backbones listed in Table 1.
@@ 2 Related Work
Domain I: Math-Reasoning
@@
Domain II: Trivia-QA
@@
Domain III: MMLU-Pro
@@
Table 4: Math correctness detection on Qwen3 1.7B-Hybrid under two max response lengths (12k, 24k). We compare Gnosis with SkyworkRM-Qwen3-8B, highlighting Gnosis’s near-constant latency and large speedups as response length increases.
@@ 3.2 Hidden-State Circuit Encoder
To summarize the sequence into a compressed representation, we then apply a Set Transformer–style encoder (lee2019set): Set Attention Blocks (SAB) followed by a Pooling-by-Multihead-Attention (PMA) block. This enables global interaction across all positions and aggregates the sequence into a small set of summary tokens, which are flattened and linearly projected to form the final hidden descriptor $z_{\text{hid}}$. Appendix A.1 details the encoder architecture and visualization, while Appendix B.3 presents full design ablations.
@@ 3.3 Attention Circuit Encoder
Because this stage operates on fixed dimensions $(L,H,d_{\text{grid}})$, the Gnosis-side compute is independent of the original sequence length $S$. Detailed architecture choices and ablations are deferred to Appendix B.2.
@@ 4.1 Experimental Setup
Training Details and Cost. We train each head for two epochs over this mixed dataset using Adam with a learning rate of $1\times 10^{-4}$. Because the backbone is frozen and all feature extractors operate at a fixed budget independent of sequence length, training is lightweight. For the largest backbone (gpt-oss-20B MoE), the full pipeline, data generation and training finishes in roughly 12 hours on $2\timesA$100 80 GB GPUs, corresponding to approximately \$25 in cloud cost. Smaller backbones train faster.
@@
Baselines. We compare against four baseline families. (1) Statistical internal scores are training-free indicators computed from the backbone’s own outputs, reported in Table 1 as Logit Entropy, Mean Token Prob, and Attn Eigenvalue Scoresriramanan2024llmcheck. (2) Trajectory/spectral internal indicators summarize cross-layer hidden-state dynamics, reported as CoE–R and CoE–Cwang2025latentcoe. (3) External judges include two open-source reward models that are state-of-the-art on public reward-model benchmarksmalik2025rewardbench, SkyworkRM-Llama3.1-8B and the family-aligned SkyworkRM-Qwen3-8Bliu2025skywork, as well as Gemini 2.5 Pro used as a judge (the Gemini judging prompt is provided in the Appendix D.1); all are reported in Table 1. (4) A Learnable probezhang2025reasoning that observes only the final answer token’s hidden state is reported separately on Qwen3 1.7B in Table 2 to isolate the limitations of single-token probing.
@@ Appendix A Architecture Overview
Figure 5: Detailed architecture of the Gnosis encoders. The Hidden Circuit Encoder (left) summarizes projected final-layer hidden states using local temporal mixing followed by global set aggregation. The Attention Circuit Encoder (right) extracts features from downsampled attention maps and models cross-layer/head structure with a lightweight axial processor. Both streams output compact, fixed-size descriptors ($z_{\text{hid}}$ and $z_{\text{attn}}$).
@@ Appendix D Backbone Outcome Statistics
Table 11: Raw outcome statistics of frozen backbones across our three evaluation domains. We report the accuracy, hallucination rate, and non-response rate for each backbone on: Math-Reasoning, Open-Domain QA (TriviaQA), and Academic Knowledge-Reasoning (MMLU-Pro). Non-response cases are reported for completeness, but are filtered out and not used in our correctness-prediction evaluation in Table 1.
Backbone Domain Accuracy Hallucination Rate Non-Response Rate

### v2

@@ 1 Introduction
Figure 1: Overview of our Gnosis self-awareness mechanism and its performance. Left: Gnosis taps hidden states and attention maps from a frozen LLM, learns to compress them into hidden/attention descriptors, and predicts a scalar correctness (hallucination) score with only $\sim$5 million extra parameters and essentially zero added inference cost. Right: Gnosis outperforms 8B Skywork reward models and a Gemini 2.5 Pro judge in AUROC on Math-Reasoning (AMC12 + AIME24/25 + HMMT Feb 2025), Open-Domain QA (TriviaQA), and Academic Knowledge (MMLU-Pro); scores are averaged over the frozen backbones listed in Table 1.
@@ 2 Related Work
Domain I: Math(AMC12 + AIME24/25 + HMMTFeb2025)
@@
Domain II: Open-Domain QA(TriviaQA)
@@
Domain III: Academic Knowledge-Reasoning(MMLU-Pro)
@@
Table 4: Correctness detection on Math-reasoning for Qwen3 1.7B-Hybrid backbone under two max response lengths (12k, 24k). We compare Gnosis with SkyworkRM-Qwen3-8B, highlighting Gnosis’s near-constant latency and large speedups as response length increases.
@@ 3.2 Hidden-State Circuit Encoder
To summarize the sequence into a compressed representation, we then apply a Set Transformer–style encoder (lee2019set): Set Attention Blocks (SAB) followed by a Pooling-by-Multihead-Attention (PMA) block. This enables global interaction across all positions and aggregates the sequence into a small set of summary tokens, which are flattened and linearly projected to form the final hidden descriptor $z_{\text{hid}}$. Figure 2 illustrates the detailed architecture design of Hidden Circuit Encoder. Appendix A.1 details the encoder architecture, while Appendix B.3 presents full design ablations.
@@ 3.3 Attention Circuit Encoder
Because this stage operates on fixed dimensions $(L,H,d_{\text{grid}})$, the Gnosis-side compute is independent of the original sequence length $S$. Figure 2 illustrates the detailed architecture design of Attention Circuit Encoder. Detailed architecture choices and ablations are deferred to Appendix B.2.
@@ 4.1 Experimental Setup
Training Details and Cost. We train each head for two epochs over this mixed dataset using Adam with a learning rate of $1\times 10^{-4}$. Because the backbone is frozen and all feature extractors operate at a fixed budget independent of sequence length, training is lightweight. For the largest backbone (gpt-oss-20B MoE), the full pipeline, data generation and training finishes in roughly 12 hours on $2\timesA$100 80 GB GPUs, corresponding to \$25 in cloud cost. Smaller backbones train faster.
@@
Baselines. We compare against four baseline families. (1) Statistical internal scores are training-free indicators computed from the backbone’s own outputs, reported in Table 1 as Logit Entropy, Mean Token Prob, and Attn Eigenvalue Scoresriramanan2024llmcheck. (2) Trajectory/spectral internal indicators summarize cross-layer hidden-state dynamics, reported as CoE–R and CoE–Cwang2025latentcoe. (3) External judges include two open-source reward models that are state-of-the-art on public reward-model benchmarksmalik2025rewardbench, SkyworkRM-Llama3.1-8B and the family-aligned SkyworkRM-Qwen3-8Bliu2025skywork, as well as Gemini 2.5 Pro used as a judge (the Gemini judging prompt is provided in the Appendix E); all are reported in Table 1. (4) A Learnable probezhang2025reasoning that observes only the final answer token’s hidden state is reported separately on Qwen3 1.7B in Table 2 to isolate the limitations of single-token probing.
@@ Appendix A Architecture Overview
Math Reasoning TriviaQA MMLU-Pro
@@ Appendix D Backbone Outcome Statistics
Table 11: Backbone outcome rates (%) across evaluation domains. For each frozen backbone, we report the fraction of instances that are correct (Accuracy), incorrect (Hallucination), or no-answer (“I don’t know”/refusal/empty). No-answer cases are shown for completeness but are filtered out from our correctness-prediction evaluation (Table 1).
Backbone Domain Backbone Accuracy Backbone Hallucination Backbone “I don’t know”